Suppose the supply of money, measured by M1, is $2.4 trillion, output, measured by real GDP, is $16.8 trillion, and the velocity of money is 6.0. Suppose the supply of money increases to $5.0 trillion but GDP and the velocity of money do not change. What is the percent by which prices change? Provide your answer as a percentage rounded to two decimal places.
Answer #1

Anwer

equation of exchange is

MV=PY

M=money supply

V=velocity

P=price level

Y=real GDP

before change P is

P=(2.4*6)/16.8

=0.857142857=0.86

after change

P=(5*6)/16.8

=1.78571429=1.79

the % change =((new P-old P)/old P)*100=((1.78571429-0.857142857)/0.857142857)*100=108.333334

=108.33%